OpenStudy (prominer10000):

The equation y = 43x represents the number of miles Alicia drives over time, where y is the number of miles and x is time in hours. The table shows the number of miles Trish drives over time. How many miles does each woman travel per hour, and who drives at a slower rate? Alicia: miles per hour Trish: miles per hour The woman who drives at a slower rate is Trish Time (h) 1.5 3.5 6 7.5 Miles 69 161 276 345

OpenStudy (prominer10000):

@563blackghost

563blackghost (563blackghost):

We are given an equation for Alicia....we want to know how many miles she drives per hour...well the equation is already the answer....we don't know of how many total hours she drives or how many hours but we do know the rate which is 43....so Alicia drives 43 miles per hour.... Now to find Trish's we would need to divide the number of miles by the time taken, we do it twice to see if it really is that unit rate.... \(\huge{\frac{69}{1.5}=?}\) \(\huge{\frac{161}{3.5}=?}\) Who drives the slowest?

OpenStudy (prominer10000):

69/1.5 = 46 and 161/3.5 equals 46 too. So Trish does 46 mph and Alicia does 43 mph, so alicia drives the slowest?

563blackghost (563blackghost):

Correct ;)